Last year, Lorat Corp. had net income of $141 million and paid out $42.3 million in the form of dividends. This year, the company has a net income of $169.2. It has identified positive NPV projects that require $152.28 million in funding. The company's target debt ratio (debt/asset) is 0.5. The company has 6 million shares outstanding.

Part 1

If the company wants to maintain the same payout ratio as last year, what should be the dividend per share (in $)?

Payout ratio last year:
d0 = D0 / NI0
= 42.3 / 141
= 0.3

Total dividends this year:
D1 = d0 * NI1
= 0.3 * 169.2
= 50.76

Dividend per share this year:
DPS1 = D1 / N
= 50.76 / 6
= 8.46 Correct ✓

Attempt 3/5 for 6 pts.

Part 2

If the company wants to maintain the same dividend per share as last year, what will be the payout ratio?
Answer #1

a). Payout Ratio (Last year) = D0 / NI0
= 42.3 / 141 = 0.3, or 30%

Total Dividends this year(D1) = Payout Ratio * NI(Current Year)

= 0.3 * 169.2 = $50.76 million

DPS(1) = D1 / Number of shares outstanding = 50.76 / 6 = $8.46

b). DPS(0) = D0 / Number of shares outstanding = 42.3 / 6 = $7.05

D1 = D0 = $42.3 million

Payout Ratio(Current Year) = D1 / NI(Current Year)

= 42.3 / 169.2 = 0.25, or 25%